# Networks with Finite VC Dimension: Pro and Contra

### v1

@@ 2 Preliminaries
In multilayer networks, units are arranged in layers, network inputs are viewed as the layer 00, output as the layer $L+1$, the layers $l=1,\dots,L$ are called hidden layers, and $L$ is called the network depths. A biologically inspired computational unit called perceptron applies a fixed activation function $\psi:{\mathbb{R}}\to{\mathbb{R}}$ to an affine functions with varying parameters. A perceptron computes functions of the form
@@
where $v\in{\mathbb{R}}^{d}$ is called a weight vector, $b\in{\mathbb{R}}$ a bias, and $v\cdot x=\sum_{i=1}^{d}v_{i}x_{i}$ is the scalar product of the weight vector with input vector $x$ (weighted sum of inputs). Originally, perceprons were endowed with sigmoidal activation functions representing hard or soft threshold, such as the Heaviside function defined as $\theta(t)=0$ for $t\ \leq 0$ and $\theta(t)=1$ for $t\geq 1$. Currently, the most widespread computational units are rectified linear units (ReLU) which are perceptrons with the activation function $\rho(t)=\max(0,t)$.
@@ 3 Probabilistic bounds on approximation and empirical errors
Let $X=\{x_{1},\dots,x_{m}\}\subset{\mathbb{R}}^{d}$, $h\in{\mathcal{B}}(X)$, :$\ell:\{-1,1\}\times\{-1,1\}\to[0,b]$ be a loss function.
@@
$\frac{1}{n}\Bigl{|}\ell(y_{1},h(x_{1}))-\ell(y^{\prime}_{1},h(x^{\prime}_{1}))\Bigr{|}\leq\frac{2b}{n}.$Box
@@ 4 Accuracy and consistency
Let $X=\{x_{1},\dots,x_{m}\}\subset{\mathbb{R}}^{d}$, ${\rm P}$ be a probability distribution on ${\mathcal{B}}(X)$ such that ${\rm P}=\prod_{i=1}^{m}{\rm P}_{i}$ where ${\rm P}_{i}$ are i.i.d., ${\mathcal{H}}\subset{\mathcal{B}}({\mathbb{R}}^{d})$, and :$\ell:\{-1,1\}\to[0,b]$ be a loss function. Then for every $h\in{\mathcal{H}}(X)$, every positive integer $n\leq m$, every random sample $S=(x_{i_{1}},y_{i_{1}}),\dots,(x_{i_{n}},y_{i_{n}})$, and every $\lambda>0$
@@ Acknowledgments
V.K. was partially supported by the Czech Science Foundation grant GA22-02067S, the institutional support of the Institute of Computer Science RVO 67985807, and a Visiting Professor grant from GNAMPA-INdAM (Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni - Instituto Nazionale di Alta Matematica). M.S. is a member of GNAMPA-INdAM and Guest Scholar at IMT - School for Advanced Studies, Lucca, Italy. He was supported by the projects PDGP DIT.AD021.104 of INM-CNR (National Research Council of Italy), where he is Research Associate, F/310027/02/X56 of the Italian Ministry of Enterprises and Made in Italy, PRIN 2022S8XSMY of the Italian Ministry of Research, and PRIN PNRR “MOTUS” (CUP: D53D23017470001), funded by the European Union – Next Generation EU program.

### v2

@@ 2 Preliminaries
In multilayer networks, units are arranged in layers, network inputs are viewed as the layer 0, output as the layer $L+1$, the layers $l=1,\dots,L$ are called hidden layers, and $L$ is called the network depths. A biologically inspired computational unit called perceptron applies a fixed activation function $\psi:{\mathbb{R}}\to{\mathbb{R}}$ to an affine functions with varying parameters. A perceptron computes functions of the form
@@
where $v\in{\mathbb{R}}^{d}$ is called a weight vector, $b\in{\mathbb{R}}$ a bias, and $v\cdot x=\sum_{i=1}^{d}v_{i}x_{i}$ is the scalar product of the weight vector with input vector $x$ (weighted sum of inputs). Originally, perceprons were endowed with sigmoidal activation functions representing hard or soft threshold, such as the Heaviside function defined as $\theta(t)=0$ for t0t\ $\leq$ 0 and $\theta(t)=1$ for $t\geq 1$. Currently, the most widespread computational units are rectified linear units (ReLU) which are perceptrons with the activation function $\rho(t)=\max(0,t)$.
@@ 3 Probabilistic bounds on approximation and empirical errors
Let $X=\{x_{1},\dots,x_{m}\}\subset{\mathbb{R}}^{d}$, $h\in{\mathcal{B}}(X)$, $\ell:\{-1,1\}\times\{-1,1\}\to[0,b]$ be a loss function.
@@
$\frac{1}{n}\Bigl|\ell(y_{1},h(x_{1}))-\ell(y^{\prime}_{1},h(x^{\prime}_{1}))\Bigr|\leq\frac{2b}{n}.$ $\Box$
@@ 4 Accuracy and consistency
Let $X=\{x_{1},\dots,x_{m}\}\subset{\mathbb{R}}^{d}$, $\rm P$} be a probability distribution on ${\mathcal{B}}(X)$ such that ${\rm P}=\prod_{i=1}^{m}{\rm P}_{i}$ where ${\rm P}_{i}$ are i.i.d., ${\mathcal{H}}\subset{\mathcal{B}}({\mathbb{R}}^{d})$, and $\ell:\{-1,1\}\to[0,b]$ be a loss function. Then for every $h\in{\mathcal{H}}(X)$, every positive integer $n\leq m$, every random sample $S=(x_{i_{1}},y_{i_{1}}),\dots,(x_{i_{n}},y_{i_{n}})$, and every $\lambda>0$
@@ Acknowledgments
V.K. was partially supported by the Czech Science Foundation grant GA25-15490S, the institutional support of the Institute of Computer Science RVO 67985807, and a Visiting Professor grant from GNAMPA-INdAM (Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni - Instituto Nazionale di Alta Matematica). M.S. is a member of GNAMPA-INdAM and Guest Scholar at IMT - School for Advanced Studies, Lucca, Italy. He was supported by the projects PDGP DIT.AD021.104 of INM-CNR (National Research Council of Italy), where he is Research Associate, F/310027/02/X56 of the Italian Ministry of Enterprises and Made in Italy, FISA-2022-00827 UAV-FIRE, PRIN 2022S8XSMY of the Italian Ministry of Research, and PRIN PNRR “MOTUS” (CUP: D53D23017470001), funded by the European Union – Next Generation EU program.